Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probabilities: Kim has 5 pairs of shoes; each pair is a diff

Expert replies
by II » Sat Aug 02, 2008 2:58 am
Kim has 5 pairs of shoes; each pair is a different color. If Kim randomly selects 2 shoes without replacement from the 10 shoes, what is the probability that she will select 2 shoes of the same color ?

(A) 2/5
(B) 1/5
(C) 1/9
(D) 1/10
(E) 1/25

I know the answer to this one ... but am interested in learning the various approaches people are taking in answering this one.

Thanks.
Join the discussion
Source: — Problem Solving |

Her first pick is inconsequential. she can choose any shoe.

For her second pick, she now has to pick the 1 shoe that matches out of the 9 remaining shoes... so 1/9. C.

This should ideally take 20-30 seconds.



II wrote:Kim has 5 pairs of shoes; each pair is a different color. If Kim randomly selects 2 shoes without replacement from the 10 shoes, what is the probability that she will select 2 shoes of the same color ?

(A) 2/5
(B) 1/5
(C) 1/9
(D) 1/10
(E) 1/25

I know the answer to this one ... but am interested in learning the various approaches people are taking in answering this one.

Thanks.
Join the discussion

by tolmar » Sat Aug 02, 2008 3:27 am
I'll go with C also
Join the discussion

by egybs » Sat Aug 02, 2008 3:30 am
You could also do :

5/10C2 = 5/45 = 1/9

but this isn't as fast as what i did before...
Join the discussion

by sudhir3127 » Sat Aug 02, 2008 5:54 am
i go with 1/9 as well...
Join the discussion

by II » Sat Aug 02, 2008 6:59 am
egybs wrote:You could also do :

5/10C2 = 5/45 = 1/9

but this isn't as fast as what i did before...
Thanks egybs.
Sorry for the basic question, buy can you please clarify what you mean by the C2 in "5/10C2"

My approach for this was:
There will be 5 pairs of shoes ... so there will be 5 colours represented. The probability of picking one colour is 2/10.
So once the first shoe is picked we have 1 more shoe which matches the one already picked ... and 9 shoes remaining ... so the probability is 1/9.
Since we are picking one shoe AND another shoe ... we multiply the two probabilities: 2/10 * 1/9 = 2/90 = 1/45

So 1/45 is the probability of picking a pair of shoes of the same colour.

Since the question doesnt specify a specific colour, and there are 5 colours ... we could pick the 1st colour OR the 2nd colour OR the 3rd colour OR the 4th colour OR the 5th colour ... (note with OR scenarios in probability, we add the probabilities).
We know the probability of picking 1 pair of shoes with the same colour is 1/45 ... so we have 1/45 + 1/45 + 1/45 + 1/45 + 1/45 = 5/45 = 1/9.

Instead of the actual answer itself, I am more interested in reviewing this problem in depth: finding out alternative approaches, key take-aways/lessons from this question, traps/tricks to be aware of etc.

Thanks.
Join the discussion

by egybs » Sat Aug 02, 2008 10:00 am
Sure,

we know that there are 5 possible ways for Kim pick the same color shoes. But how many different ways can she select them?

10C2 ("10 choose 2") is the number of ways that you can choose 2 things out of 10 things.

The operation works as follows:

10!/(2!*(10-2)!)

This simplifies to:
10*9/2 = 45.

So Kim could choose the two shoes 45 different ways. 5/45 = 1/9.

Your method, while correct, is unnecessarily long. It doesn't matter what colour shoe she chooses first. Regardless of what color shoe she chooses first, she'll need to pick that same color shoe again out of the nine remaining shoes. which is 1/9.

This is very similar to the 4 letters in envelopes question. Where it doesn't matter which envelope is correctly placed:
https://www.beatthegmat.com/probability- ... 14363.html

In that question you could either do 4* 1/4 * 2/3 * 1/3 or you could just do 2/3 *1/3

Hope this helps...

II wrote:
egybs wrote:You could also do :

5/10C2 = 5/45 = 1/9

but this isn't as fast as what i did before...
Thanks egybs.
Sorry for the basic question, buy can you please clarify what you mean by the C2 in "5/10C2"

My approach for this was:
There will be 5 pairs of shoes ... so there will be 5 colours represented. The probability of picking one colour is 2/10.
So once the first shoe is picked we have 1 more shoe which matches the one already picked ... and 9 shoes remaining ... so the probability is 1/9.
Since we are picking one shoe AND another shoe ... we multiply the two probabilities: 2/10 * 1/9 = 2/90 = 1/45

So 1/45 is the probability of picking a pair of shoes of the same colour.

Since the question doesnt specify a specific colour, and there are 5 colours ... we could pick the 1st colour OR the 2nd colour OR the 3rd colour OR the 4th colour OR the 5th colour ... (note with OR scenarios in probability, we add the probabilities).
We know the probability of picking 1 pair of shoes with the same colour is 1/45 ... so we have 1/45 + 1/45 + 1/45 + 1/45 + 1/45 = 5/45 = 1/9.

Instead of the actual answer itself, I am more interested in reviewing this problem in depth: finding out alternative approaches, key take-aways/lessons from this question, traps/tricks to be aware of etc.

Thanks.
Join the discussion